# ltxutilRepair the float

### v1

@@ Abstract
Connecting multiple processors via quantum interconnect technologies could help to overcome issues of scalability in single-processor quantum computers. Transmission via these interconnects can be performed more efficiently using quantum multiplexing, where information is encoded in high-dimensional photonic degrees of freedom. We explore the effects of multiplexing on logical error rates in surface codes and hypergraph product codes. We show that, although multiplexing makes loss errors more damaging, assigning qubits to photons in an intelligent manner can minimize these effects, and the ability to encode higher-distance codes in a smaller number of photons can result in overall lower logical error rates. This multiplexing technique can also be adapted to quantum communication and multimode quantum memory with high-dimensional qudit systems.
@@ 1 Introduction
Quantum computers are expected to solve problems that are intractable using classical computation [1, 2], but these powerful quantum algorithms require high qubit counts and deep circuits to solve problems of interesting size [3, 4]. While the qubit counts of quantum processors have been increasing rapidly in recent years, various physical constraints impose limits on the possible size of a single quantum processor [5, 6]. Quantum interconnects provide a resolution to this problem by allowing for the networking and cooperative operation of multiple quantum processors [7], as well as the use of separate quantum memories [8], quantum repeaters [9, 10, 11, 12] & networks [13, 14, 15] in analogy with classical computing architectures. Optical systems are considered leading candidates for practical implementations of quantum interconnects [16] and also as quantum memories [17] due to long coherence times [18].
@@
Due to the high noise levels inherent in quantum systems, large-scale quantum algorithms cannot be executed reliably without the use of quantum error correcting codes (QECCs) [19, 20], which enable fault-tolerant quantum computation (FTQC) [21, 22]. Similarly, optical interconnects [7, 16] can suffer from high photon loss rates and so QECCs should be used to protect information transmitted through these channels. In principle, it may be preferable to use different codes for these different settings [23], but in practice, the transfer of information between these different codes may be challenging enough that it is easier to use only a single code. For instance, fault-tolerant logic with surface codes [24] has been very well studied [25, 26], while quantum Reed-Solomon codes [27] provide efficient and loss-tolerant protection for transmission through optical channels, but it is not clear how to interface or switch between these two families codes. Therefore, for performing distributed computation with interconnects, using surface codes (or their generalizations) for both computation and transmission [28] is a natural alternative to using multiple codes. This is much less efficient in the sense of error-correction capability for the communication part, but these overheads can be reduced using quantum multiplexing [29].
Figure 1: An example of an optical circuit encoding $2^{2}$-dimensional quantum information into a single photon.
@@
Quantum multiplexing is a technique for encoding high-dimensional quantum information onto a single photon by exploiting multiple different photonic degrees of freedom (DoF) or a single multi-component degree of freedom. Such encodings can be performed using only linear optical elements and can significantly reduce the resources associated with quantum communication [30, 31, 32]. In this work, we examine the potential of quantum multiplexing for enabling efficient transmission of surface and hypergraph product (HGP) codes through optical channels. Densely encoding many qubits of these codes into small numbers of photons has the potential to make loss errors much more damaging, but we present various techniques (e.g. optimized strategies for qubit-to-photon assignment) that can mostly or completely eliminate these downsides.
@@
The rest of this paper is organized as follows. In Sec. 2 we review the relevant background for quantum multiplexing and communication over lossy optical channels. Then in Sec. 3 we propose three approaches to error-corrected quantum communication using multiplexing that provide different ways of reducing the impact of loss errors. This is followed up in Sec. 4 and Sec. 5 with an examination of some of these approaches in more detail for surface codes and HGP codes respectively. Finally, we discuss our findings and conclude this work in Sec. 6.
@@ 2 Background
In this section, we briefly overview quantum multiplexing and erasure correction and show how these elements appear in practical quantum communication protocols.
@@ 2.1 Quantum Multiplexing
In photon-based quantum information processing, various degrees of freedom (DOF) can be utilized to encode qubits. Polarizations [33], time-bins [34, 35, 36], paths (dual rail) [37], orbital angular momentum [38], and frequency-bin [39, 40] are typical examples of DOF in a single photon which are commonly used in experiment. Multi-level time-bins make it especially easy to encode high-dimensional quantum information in a single photon. For instance, Fig. 1 shows a method for encoding higher dimensional information ($2^{2}$-dimension) using polarization and time-bin DOF. This circuit takes a photon whose polarization is encoded with quantum information as input. This input photon has one qubit of information. After passing through this circuit, the photon has both polarization and time-bin degrees of freedom. The polarization encodes a two-dimensional Hilbert space, and the time-bin encodes a four-dimensional one. Therefore, the Hilbert space of the final encoded state has dimension $4$ encoding, thus, $2$ qubits of information. This encoding can easily generalized to higher dimensional multiplexed photons as shown in Appendix A. Significantly, encoding high-level time-bin states only requires linear optical elements and classical optical switches.
Quantum multiplexing [29] is a method to encode higher dimensional quantum information in a single photon using these multiple degrees of freedom. In this work, we consider encoding $2^{m}$-dimensional quantum information using $m$ components of a DOF per photon where $m$ is an integer ($m=1$ corresponds to no multiplexing).
@@ 2.2 Erasure Channel and Correction
After replacing the qubits, one can perform stabilizer measurements as normally occurs in surface codes. Then, the erasure is converted into random Pauli errors with the exact probabilities ($1/4$) for $\{I,X,Y,Z\}$. This random Pauli can also be regarded as independent $X$ and $Z$ errors with a probability of $1/2$. This allows for the decoding of an erasure error. The (surface code) peeling decoder [55], which is a linear-complexity erasure decoder using this procedure, has been proposed as a maximum-likelihood decoder for erasure errors in the surface code. We briefly overview the peeling decoder and its surface code generalization in Appendix C.
2.3 Applying quantum multiplexing to error-corrected erasure channel
We describe the steps to perform error-corrected quantum communication over a multiplexed erasure channel as an example with surface codes illustrated in Fig. 2. As the first step, the sender prepares an encoded quantum state. Then, in the second step, the sender assigns and converts each physical data qubit to a photon. In the conventional case, different qubits are assigned to different photons, whereas when quantum multiplexing is in use, different qubits can be assigned to the same photon. In the instance of Fig. 2, qubits 00 and $6$ are attached to photon 00, qubits $1$ and $7$ are attached to photon $1$, etc. We will discuss the optimal assignment strategy later. For the third step, the codeword then goes through an optical channel, which has a loss error. During the transmission, some photons can be lost, causing the loss of all the qubits attached to them as well. For instance, when photon 1 is lost, qubits 1 and 7 will also be lost, as shown in Fig. 2. In the fourth step, the receiver reconstructs (imperfect) codewords with remaining qubits. As the final step, the receiver converts the qubit loss to random unitary as explained in Sec. 2.2. Then, the decoding algorithm estimates the errors, and the receiver performs correction. In the second part of the paper, these steps were simulated to obtain the performance of communication.
@@ 2.3 Applying quantum multiplexing to error-corrected erasure channel
Figure 2: Flow of quantum communication with surface code using multiplexed photons. In the first step, a quantum state is encoded into a surface code. Each circle with a number inside is the physical data qubit, and the grey circles without any number are auxiliary qubits used for stabilizer measurement. For the second step, in a quantum multiplexing scenario, one assigns each physical qubit of the codeword to single photons using an assignment strategy. For instance, in this figure, two components of the time-bin DOF in each photon are used so that each photon can encode two qubits. There is a degree of freedom in which qubit is assigned to which photon, so it is required to make a map function. We call this function the interleaving assignment strategy. Here, the colors of the qubits indicate which photon the qubit is encoded to, which is the result of the assignment strategy. Then, the encoded photons go over a lossy channel. Here, we assume that we know which photons have been lost during the transmission (erasure channel). If a photon has been lost, all the qubits in the photon have been lost. Finally, we demultiplex and decode it to a code word of the surface code using the peeling decoder [55] and a correction method for erasure error shown in Sec. 2.2.
@@ 3 Multiplexed quantum communication with error-correcting codes
We propose three different scenarios in which multiplexing is used to enhance the efficiency of quantum communication. In each case, $m$ qubits are encoded into each photon, and we compare them to the case of transmitting a codeword of a given code $C$ without multiplexing ($m=1$). ). The three scenarios are
@@
    $m$ codewords of $C$ are transmitted using the same number of photons as the $m=1$ case.
@@
    An $m$-times larger code from the same family as $C$ is transmitted using the same number of photons as the $m=1$ case.
@@
Table 1: Comparison of the surface code communication without multiplexing and three scenarios with multiplexing. Parameters that are improved by multiplexing are in red fonts. The case without multiplexing requires one qubit per photon. (A) The first scenario is only applicable when sending multiple codewords. This enables one to send more codewords with the same number of photons, drastically improving the channel’s throughput. (B) The second scenario sends the same number of codewords with the bigger code, improving the error tolerance. (C) The third scenario sends the same codeword with fewer photons, drastically improving the channel’s throughput. The number of photons required in scenario (C) is $\lfloor 2d^{2}/m\rfloor$, where $\lfloor x\rfloor$ is the floor function of $x$.
@@
(B) Sending $m$times bigger codewords
In the second scenario, a larger number of qubits are used to encode a single codeword from a code in the same family with an $m$-fold longer length. If this scenario is applied to the surface code, it achieves $\sqrt{m}$ times larger distance than the no-multiplexing case ($m=1$).
Fig. 3 shows a Monte Carlo simulation of the logical $Z$ error rate for this scenario for the surface code. To determine whether a logical $Z$ error occurred, we checked whether the errors left after the decoding process were anti-commutative with a logical $X$ operator of any logical qubit in the codeword. Each data point in the simulation is obtained from $10^{5}$ shots, and the error bar is given by the Agresti–Coull interval [56]. This scenario introduces correlations in errors between the qubits in the code, which may degrade the performance. However, if $m$ is sufficiently small relative to the code size, the benefit gained by increasing the code size is more significant. All the programs we used to simulate multiplexed quantum communication with surface codes are available here [57]. The logical error rate significantly decreases as the code size and $m$ increase.
Note that the logical $Z$ error rate converges to $0.75$. This is because there are two logical qubits in the codes, and the logical $Z$ error for each qubit converges to $0.5$; hence, the probability that both qubits are logical $Z$ error-free is $0.25$. In practice, this does not mean the encoded information is recovered, though.
Figure 3: Performance of $[\![2d^{2},2,d]\!]$ toric codes in scenario (B) with about $100$ photons. Each curve shows the case with different code sizes and the number of qubits encoded in each photon. The logical error rate can be reduced by increasing the number of qubits per photon $m$ and the code distance $d$.
@@
In the third scenario, a smaller number of photons are used to encode a single codeword. The code parameters are the same as the case without multiplexing. It has no restriction on the number of codewords and can improve the efficiency of surface code communication in general. This method introduces a correlation to the errors. Fig. 4 shows this scenario’s logical $Z$ error rate versus the photon loss probability for different values of $m$. It shows that as $m$ increases, the logical error rate decreases.
Figure 4: Scenario (C) multiplexing performance for $[\![200,2,10]\!]$ toric code with multiplexing using different values of $m$ (the number of qubits per photon). The assignment of qubits to photons is uniformly random. Increasing $m$ allows code words to be transmitted with fewer photons, but the logical error rate increases because multiple qubits in the same photon have strongly correlated errors.
While the number of photons is less compared to the no-multiplexing case, the effects of the correlated errors can be very detrimental to the performance of such a system. A more suitable assignment of the qubits to the multiplexed photos can ideally reduce those detrimental effects. In the next subsection, we explore five different strategies of qubits assignment.

### v2

@@ Abstract
Connecting multiple processors via quantum interconnect technologies could help overcome scalability issues in single-processor quantum computers. Transmission via these interconnects can be performed more efficiently using quantum multiplexing, where information is encoded in high-dimensional photonic degrees of freedom. We explore the effects of multiplexing on logical error rates in surface codes and hypergraph product codes. We show that, although multiplexing makes loss errors more damaging, assigning qubits to photons in an intelligent manner can minimize these effects, and the ability to encode higher-distance codes in a smaller number of photons can result in overall lower logical error rates. This multiplexing technique can also be adapted to quantum communication and multimode quantum memory with high-dimensional qudit systems.
@@ 1 Introduction
Quantum computers are expected to solve problems that are intractable using classical computation [1, 2], but these powerful quantum algorithms require high qubit counts and deep circuits to solve problems of interesting size [3, 4]. While the qubit counts of quantum processors have been increasing rapidly in recent years, various physical constraints impose limits on the possible size of a single quantum processor [5, 6]. Quantum interconnects provide a resolution to this problem by allowing for the networking and cooperative operation of multiple quantum processors [7], as well as the use of separate quantum memories [8], quantum repeaters [9, 10, 11, 12] and networks [13, 14, 15] in analogy with classical computing architectures. Optical systems are considered leading candidates for practical implementation of quantum interconnects [16] and also as quantum memories [17] due to their long coherence times [18].
@@
Due to the high noise levels inherent in quantum systems, large-scale quantum algorithms cannot be executed reliably without the use of quantum error correcting codes (QECCs) [19, 20], which enable fault-tolerant quantum computation (FTQC) [21, 22]. Similarly, optical interconnects [7, 16] can suffer from high photon loss rates and so QECCs should be used to protect information transmitted through these channels. In principle, it may be preferable to use different codes for these different settings [23], but in practice, the transfer of information between these different codes may be challenging enough that it is easier to use only a single code. For instance, fault-tolerant logic with surface codes [24] has been very well studied [25, 26], while quantum Reed-Solomon codes [27] provide efficient and loss-tolerant protection for transmission through optical channels, but it is not clear how to interface or switch between these two families of codes. Therefore, for performing distributed computation with interconnects, using surface codes (or their generalizations) for both computation and transmission [28] is a natural alternative to using multiple codes. This is much less efficient in the sense of error-correction capability for the communication part, but these overheads can be reduced using quantum multiplexing [29].
[…]
@@
Quantum multiplexing is a technique for encoding high-dimensional quantum information into a single photon by exploiting multiple different photonic degrees of freedom (DOF) or a single multi-component degree of freedom. Such encodings can be performed using only linear optical elements and can significantly reduce the resources associated with quantum communication [30, 31, 32]. In this work, we examine the potential of quantum multiplexing for enabling efficient transmission of surface and hypergraph product (HGP) codes through optical channels. Densely encoding many of these codes’ qubits into small numbers of photons has the potential to make loss errors much more damaging, but we present various techniques (e.g. optimized strategies for qubit-to-photon assignment) that can mostly or completely eliminate these downsides.
@@
The rest of this paper is organized as follows. In Sec. 2 we review the relevant background for quantum multiplexing and communication over lossy optical channels. Then in Sec. 3 we propose three approaches to error-corrected quantum communication using multiplexing that provide different ways of reducing the impact of loss errors. This is followed up in Sec. 4 and Sec. 5 with an examination of some of these approaches in more detail for surface codes and HGP codes, respectively. Finally, we discuss our findings and conclude this work in Sec. 6.
@@ 2 Background
In this section, we provide a brief overview of quantum multiplexing and erasure correction and show how these elements appear in practical quantum communication protocols.
@@ 2.1 Quantum Multiplexing
In photon-based quantum information processing, various DOFs can be utilized to encode qubits. Polarizations [33], time-bins [34, 35, 36], paths (dual rail) [37], orbital angular momentum [38], and frequency-bin [39, 40] are typical examples of DOFs in a single photon which are commonly used in experiments. Multi-level time-bins make it especially easy to encode high-dimensional quantum information in a single photon. For instance, Fig. 1 (A) shows a method for encoding higher dimensional information (dimension $2^{2}$ for two qubits) using polarization and time-bin DOF. The series of Half-Quarter-Half waveplates as shown in Fig. 1 (A) as HQH allows one to apply an arbitrary unitary gate to the polarization DOF of a photon. This corresponds to applying rotation gates around a Z-X-Z basis in a qubit. We label the horizontal and vertical components of a polarized photon with $H$ and $V$, respectively, and the short and long components of a temporal mode (time-bin) with $S$ and $L$, respectively.
[…]
The polarization encodes a 2-dimensional Hilbert space, and the time-bin encodes another 2-dimensional Hilbert space. Therefore, the Hilbert space of the final encoded state has dimension $4$, thus encoding $2$ qubits of information. This encoding can easily be generalized to higher dimensional multiplexed photons as shown in Appendix B. Significantly, encoding high-level time-bin states only requires linear optical elements and classical optical switches.
Quantum multiplexing [29] is a method to encode higher dimensional quantum information into a single photon using these multiple degrees of freedom. One can transfer the state of this multiplexed photon embedded with an extra DOF (time-bin) into two matter qubits [29] as shown in Fig. 1 (B). First, the photon is entangled with the internal DOFs of a matter qubit, depicted as a red dot in a box in the Figure. Then the $HL$ and $VS$ components of the photon are switched by the switching setup module (Sw). This is a slightly modified setup in [29] where we use conventional PBSs rather than 45°-PBSs. Next, the photon interacts with another matter qubit. Here we label the basis states of a matter qubit |$\ket{0}$ and |$\ket{1}$. After measuring the DOFs of the photon, the state of the photon is projected into the matter qubits upon a Pauli frame correction depending on the measurement outcome (write-in). This procedure can be extended to photons with an arbitrary number of time-bin components by adding the corresponding number of matter qubits. We can recover the information stored in the matter qubits with the process shown in Fig. 1 (C) (read-out). This process can be regarded as the inverse of Fig. 1 (B). Here, a polarized photon interacts with a matter qubit $m_{2}$. Next, the photon goes through the Sw module which switches the $HL$ component and the $VS$ component. Then, the photon interacts with another matter qubit $m_{1}$. Finally, both matter qubits are measured in the $X$-basis. This projects the state of the matter qubits into the multiplexed photon. We give a quantum circuit representation in Fig. 1 (D), and the full description of the state transition in the Appendix A for Fig. 1 (B) and (C).
In this work, we consider encoding $2^{m}$-dimensional quantum information using $m$ components of a DOF per photon where $m$ is an integer ($m=1$ corresponds to no multiplexing).
@@ 2.2 Erasure Channel and Correction
After replacing the qubits, one can perform stabilizer measurements as normally occurs in surface codes. Then, the erasure is converted into random Pauli errors with the exact probabilities ($1/4$) for $\{I,X,Y,Z\}$. This random Pauli can also be regarded as independent $X$ and $Z$ errors with a probability of $1/2$. This allows for the decoding of an erasure error. The (surface code) peeling decoder [55], which is a linear-complexity erasure decoder using this procedure, has been proposed as a maximum-likelihood decoder for erasure errors in the surface code. We briefly overview the peeling decoder and its surface code generalization in Appendix D.
2.3 Applying quantum multiplexing to an error-corrected erasure channel
@@
Figure 2: Flow of quantum communication with the surface code using multiplexed photons. In the first step, a quantum state is encoded into a surface code. Each circle with a number inside denotes the physical data qubits, and the grey circles without any numbers are auxiliary qubits used for stabilizer measurements. For the second step in a quantum multiplexing scenario, one assigns each physical qubit of the codeword to single photons using an assignment strategy. For instance, in this figure, two components of the time-bin DOF in each photon are used so that each photon can encode two qubits. There is freedom in which qubit is assigned to which photon, so it is necessary to define a function mapping qubits to photons. We call this function the assignment strategy. Here, the colors of the qubits indicate the photon to which it is assigned based on the assignment strategy. Then, the encoded photons pass through a lossy channel. Here, we assume that we know which photons have been lost during the transmission (erasure channel). If a photon has been lost, all the qubits assigned to this photon have been lost. Finally, we demultiplex the received photons and decode it to a codeword of the surface code using the peeling decoder [55] and a correction method for erasure errors as described in Sec. 2.2.
We describe the steps to perform error-corrected quantum communication over a multiplexed erasure channel using the surface code as an example, as illustrated in Fig. 2. For the first step, the sender prepares an encoded quantum state. Then, in the second step, the sender assigns and converts physical data qubits to photons. In the conventional case, different qubits are assigned to different photons, whereas when using quantum multiplexing, different qubits can be assigned to the same photon. In the instance of Fig. 2, qubits 00 and $6$ are assigned to photon 00, qubits $1$ and $7$ are assigned to photon $1$, etc. We will discuss the optimal assignment strategy later.
For the third step, a codeword passes through an optical channel which has the possibility of erasure errors. During the transmission, some photons can be lost, causing the loss of all qubits assigned to these photons as well. For instance, when photon $1$ is lost, qubits $1$ and $7$ will also be lost as shown in Fig. 2, resulting in an erasure error. In the fourth step, the receiver reconstructs an incomplete codeword missing some qubits from the non-erased qubits assigned to the remaining photons. For the final step, the receiver replaces the missing qubits with mixed states as explained in Sec. 2.2. Effectively, this results in a codeword with random Pauli errors on some qubits. Next, the decoding algorithm estimates the errors, and the receiver performs a correction. In the second half of this paper, we examine the performance of this communication procedure via numerical simulation.
@@ 3 Multiplexed quantum communication with error-correcting codes
We propose three different scenarios in which multiplexing is used to enhance the efficiency of quantum communication. In each case, $m$ qubits are encoded into each photon, and we compare them to the case of transmitting a codeword of a given code $C$ without multiplexing ($m=1$). The three scenarios are:
@@
    $m$ codewords of $C$ are transmitted using the same number of photons as the $m=1$ case;
@@
    an $m$-times larger code from the same family as $C$ is transmitted using the same number of photons as the $m=1$ case;
@@
Table 1: Comparison of surface code communication without multiplexing and three scenarios with multiplexing. Red fonts show parameters that are improved by multiplexing. The case without multiplexing requires one qubit per photon. (A) The first scenario is only applicable when sending multiple codewords. This enables one to send more codewords with the same number of photons, drastically improving the channel’s throughput. (B) The second scenario sends the same number of codewords but uses a larger code, thus improving the error tolerance. (C) The third scenario sends the same codeword with fewer photons, which also significantly improves the channel’s throughput. The number of photons required in scenario (C) is $\lfloor 2d^{2}/m\rfloor$, where $\lfloor x\rfloor$ is the floor function of $x$.
@@
(B) Sending $m$-times bigger codewords
In the second scenario, a larger number of qubits is used to encode a single codeword from a code in the same family with an $m$-fold longer length. If this scenario is applied to the surface code, it achieves $\sqrt{m}$ times larger distance than the no-multiplexing case ($m=1$).
[…]
Readers may wonder why the logical $Z$ error rate converges to $0.75$, even though for a loss probability of $1$ the code will be completely erased. This is due to the fact that, following such a complete erasure, the recovery procedure will reinitialize the code in a random state subject to one of the logical operators $II,ZI,IZ,ZZ$ with equal probability, and so $25\%$ of the time we do not observe a logical $Z$ error. Importantly, this finite recovery probability is observed because, in this case, we care only about the presence/absence of logical $Z$ operators, and it does not imply that an arbitrary quantum state can be recovered.
Figure 3: Performance of $[\![2d^{2},2,d]\!]$ surface codes in scenario (B) with about $100$ photons. Each curve shows the case with different code sizes and the number of qubits encoded in each photon. The logical error rate can be reduced by increasing the number of qubits per photon $m$ and the code distance $d$.
@@
In the third scenario, a smaller number of photons are used to encode a single codeword. The code parameters are the same as the case without multiplexing. It has no restriction on the number of codewords and can improve the efficiency of surface code communication in general. This method introduces a correlation to the errors. Fig. 4 shows this scenario’s logical $Z$ error rate versus the photon loss probability for different values of $m$. It shows that as $m$ increases, the logical error rate also increases.
Figure 4: Scenario (C) multiplexing performance for the $[\![200,2,10]\!]$ surface code with multiplexing using different values of $m$ (the number of qubits per photon). The assignment of qubits to photons is uniformly random. Increasing $m$ allows code words to be transmitted with fewer photons, but the logical error rate increases because multiple qubits in the same photon have strongly correlated errors.
Although the number of photons is reduced compared to the no-multiplexing case, the effects of the correlated errors can be very detrimental to the performance of such a system. Ideally, these detrimental effects can be reduced by strategically assigning qubits to the multiplexed photons. In the next subsection, we explore five different strategies for qubit assignment.
3.1 Limitation and fault-tolerance of multiplexed communication
Generating time-bin multiplexed photons requires linear optical elements such as polarizing beam splitters, delay lines, wave plates, and optical switches, as shown in Fig. 1 (A). In particular, the latter suffer from loss errors that are detrimental to the overall performance of the system. In [29], the authors analyze the impact of imperfect optical switches on the performance of a purification protocol. They show that, for an optical switch efficiency equal to 99%, adding one time-bin mode to a polarized photon only slightly reduces the performance of the system. However, when several optical switches are in use to encode multiple time-bin modes, much higher efficiencies are required. Although not fully achieved with the current devices, future high-efficiency optical switches are required for several quantum technologies.
While it is possible for multiplexing to lead to correlated errors between qubits stored in the same photon, we do not expect this to compromise fault tolerance except in the case where these qubits correspond to the support of a logical operator of the code, and the majority of our assignment strategies are designed to prevent this. More generally, if we require that the number of qubits per photon is a) independent of the size of the code and b) smaller than the code distance, then these correlated errors can, at most, result in a constant-factor reduction of the distance, and so long as the non-multiplexed procedure is fault-tolerant, the multiplexed procedure will be fault-tolerant as well.